### v1

# SonicID: User Identification on Smart Glasses with Acoustic Sensing

Abstract.
Smart glasses have become more prevalent as they provide an increasing number of applications for users. They store various types of private information or can access it via connections established with other devices. Therefore, there is a growing need for user identification on smart glasses. In this paper, we introduce a low-power and minimally-obtrusive system called SonicID, designed to authenticate users on glasses. SonicID extracts unique biometric information from users by scanning their faces with ultrasonic waves and utilizes this information to distinguish between different users, powered by a customized binary classifier with the ResNet-18 architecture. SonicID can authenticate users within 0.12 seconds, with an energy consumption of 19.8 mAs per trial. A user study involving 24 participants confirms that SonicID achieves a true positive rate of 96.5%, a false positive rate of 4.1%, and a balanced accuracy of 96.2% using just 4 minutes of training data collected for each new user. This performance is relatively consistent across different remounting sessions and days. Given this promising performance, we further discuss the potential applications of SonicID and methods to improve its performance in the future.
User Identification, Smart Glasses, Acoustic Sensing, Machine Learning
copyright: nonecopyright: acmcopyrightjournal: IMWUTjournalyear: 2024journalvolume: 0journalnumber: 0publicationmonth: 0price: 15.00doi: xxxxxx/xxxxxxx
1. Introduction
Smart glasses, such as Vuzix Smart Glasses (Vuzix, 2024), Lenovo ThinkReality A3 (Lenovo, 2024), Meta Ray-Ban (Meta, 2024), Snap Spectacles 3 (Snap, 2024), etc., have been growing in popularity and providing users with more and more interactive applications in recent years. Similar to smartphones and other wearable devices (e.g. smart watches and VR headsets), smart glasses start to store users’ personal data and private information including their private messages, photos, videos and banks accounts. Leakage of private information and inconvenience might be induced to certain users if bad actors obtain this information from their smart glasses. Moreover, some smart glasses are connected to other devices such as smartphones and smart TVs. It is possible for smart glasses without user authentication to be utilized by bad actors to give unwanted commands to or display inappropriate contents on these connected devices. Therefore, there is an increasing need for implementing user authentication methods on smart glasses.
However, it is difficult to directly apply traditional frontal-camera-based authentication techniques, e.g. FaceID, on glasses because they require placing a camera in front of the user to capture their face. Researchers have put efforts into deploying cameras on glasses to capture biometric information of the user from their facial contours (Lim et al., 2023), eye movements (Zhang et al., 2018), or iris features (Li and Po-Jen, 2017). Though these camera-based methods achieve satisfactory performance of user authentication on glasses, they require instrumenting a camera which is relatively large compared to the small size of glasses and might partially block the user’s view. Furthermore, the cameras on commodity smart glasses, though can be smaller in size, usually point forwards to capture the environmental information. Barriers exist if they are to be directly used to capture the biometric information of the user’s face. Password-based authentication methods have been explored on smart glasses as well, especially with voice input (Bailey et al., 2014; Yadav et al., 2015; Li et al., 2021). In addition, behavioral biometrics, such as tapping gestures (Islam et al., 2018), touch gestures (Jagmohan et al., 2016; Peng et al., 2017), and voice commands (Peng et al., 2017), have been leveraged to authenticate users. These methods require the user’s interaction and are notably vulnerable to shoulder-surfing attacks since these gestures are entered on a touchpad and the vocalized commands can be overheard by bystanders.
In this paper, we present SonicID, a low-power and minimally-obtrusive solution to user authentication on smart glasses, which identifies users automatically as soon as they put on the glasses, using active acoustic sensing. We place one pair of speaker and microphone on each side of a pair of glasses. The sensors are instrumented on the hinges of the glasses, pointing downwards at the user’s face. The core sensing principle of SonicID is to use a sonar-like technique to scan the shape of the user’s face as the biometric information for authentication. Similar to how FaceID scans the user’s face with infrared cameras, when a user wears the glasses instrumented with SonicID, it scans the user’s face automatically with acoustic signals, and then compares scanned patterns with the profiles of registered users to determine whether or not this user can have access to the device.
Specifically, the speakers on the glasses emit encoded inaudible Frequency Modulated Continuous Waves (FMCW) towards the user’s face after they put on the glasses. The signals continuously and repeatedly sweep at different frequency ranges on two sides to avoid interference. The transmitted signals are reflected by the user’s face and captured by the microphones on the glasses. After signal processing using the received signals and transmitted signals, SonicID obtains unique acoustic patterns related to this user’s face. Considering everyone has different shapes of faces, the user’s biometric information is included in the captured acoustic patterns which can be learned to distinguish this user from other users and thus authenticate them securely. In SonicID, we feed the processed acoustic patterns into a customized binary classifier using the ResNet-18 architecture to extract distinct features so as to authenticate users. The classifier is trained with the training data collected when a user registers on the smart glasses and it keeps authenticating users in later usage of the device just like traditional user authentication techniques.
To validate the performance of our SonicID system, we conducted a user study with 24 participants. The study results demonstrate that a new user needs to provide 4 minutes of data to training the binary classifier on the first day when they use the smart glasses, to achieve a true positive rate of 96.5%, a false positive rate of 4.1% and a balanced accuracy of 96.2%. In this study, we also asked all participants to come back to test the performance of the system on two other different days. The results show that the performance of the system remains relatively consistent across three different days, especially for the false positive rate. If we continue to reduce the amount of training data to 1 minute for each participant, a true positive rate of 90.3%, a false positive rate of 1.6% and a balanced accuracy of 94.3% are achieved on Day 1. This study validates that SonicID achieves a satisfactory and stable authentication performance on smart glasses. Compared with frontal-camera-based authentication techniques such as FaceID, SonicID does not need a camera to be placed in front of the user. Other traditional authentication methods which take fingerprints or passwords as input can eliminate the need for a frontal camera. However, they require users’ operations for authentication while SonicID automatically logs valid users in when they put the device on. Moreover, SonicID is capable of successfully authenticating users within 0.12 seconds with a energy consumption of 19.8 mAs every time it authenticates the user. This low-power feature guarantees that SonicID can work approximately 28,000 times theoretically with the battery of Meta Ray-Ban (154 mAh) if no other operations are running on the device.
The main contributions of this paper are summarized as follows:
  • We proposed a low-power and minimally-obtrusive solution to the user authentication task on smart glasses, using active acoustic sensing. It demonstrated that the shape of the user’s face can be scanned with acoustic signals as biometric information for authentication.
  • We prototyped the system with one pair of speaker and microphone on each side of the glasses. The system maintains low-power, consuming only 19.8 mAs of energy every time it authenticates the user.
  • We conducted a user study with 24 participants, validating that only 4 minutes of training data for each new user guarantees that the system achieves a true positive rate of 96.5%, a false positive rate of 4.1% and a balanced accuracy of 96.2%. The performance remains relatively consistent in evaluations across different days.
2. Related Work
In this section, we present the related work of SonicID in the following three scopes: (1) Authentication methods on non-wearables; (2) Authentication methods on wearable devices; (3) Comparison between our work and other authentication methods on wearables.
2.1. Authentication on Non-Wearable Devices
Authentication mechanisms are critical in devices like smartphones, tablets, and computers that store sensitive personal information, provide access to banking services, and control security systems (Karlson et al., 2009). Traditional authentication methods, such as passwords and Personal Identification Numbers (PINs), are prevalent but not without shortcomings. Users may forget these credentials, and they are susceptible to security breaches, including shoulder-surfing attacks (Tari et al., 2006). An alternative approach, pattern-based passwords, leverages the pictorial superiority effect by allowing users to remember shapes, offering a more intuitive recall process (Jermyn et al., 1999; Nelson et al., 1976; Standing, 1973). However, these patterns are vulnerable to both shoulder-surfing and smudge attacks (Aviv et al., 2010).
Biometric authentication has gained popularity due to its perceived security and ease of use. Fingerprint sensing is widely used in current devices (Support, [n.d.]b). Face recognition technology, although increasingly common, can be compromised using photographs or videos of the user (Duc and Minh, 2009). To counter this, systems like Apple’s Face ID employ 3D reconstruction of facial features (Support, [n.d.]a). Iris scans offer another biometric alternative, noted for their uniqueness and difficulty to replicate (Daugman, 1994). Additionally, multimodal biometric systems have been explored, combining different physical attributes for enhanced security. For instance, Rokita et al. (Rokita et al., 2008) integrates face and hand images, while Marcel et al. (Marcel et al., 2010) combines face and speech recognition. EchoPrint (Zhou et al., 2018) merges facial recognition with acoustic analysis to counter spoofing attempts. VoiceLive (Zhang et al., 2016) and VoiceGesture (Zhang et al., 2017) utilize unique vocal characteristics and articulatory gestures, respectively, to thwart replay attacks. AirSign (Shao et al., 2020) innovatively employs acoustic sensing and motion sensors for air signature-based authentication.
Behavioral biometrics also provide promising avenues for authentication. These methods analyze user-specific behaviors, such as pattern input dynamics (De Luca et al., 2012), activity patterns (ul Haq et al., 2017; Malik et al., 2019), gait (Zou et al., 2020), and even unique hand movements and orientations (Sitová et al., 2016). BreathPrint (Chauhan et al., 2017) further extends this domain by using audio features from individual breathing patterns. Such techniques offer continuous and passive authentication, enhancing both security and user convenience.
However, these mature authentication technologies above usually cannot be easily transplanted to wearable devices considering their small size and limited battery capacity.
Table 1. SonicID and Other Authentication Methods on Wearables. Some papers reported Equal Error Rate (EER) which was converted to Balanced Accuracy ($BAC=1-EER$) in this table. The balanced accuracy in this table is the cross-session performance if the system was evaluated across different remounting sessions.
Project Form Sensors Biometrics User Cross- Cross- Study Balanced
Factor Extracted Interaction? Session? Day? Participants Accuracy
SonicID Glasses Acoustics Face No Yes Yes 24 96.2%
Isobe et al. (Isobe and Murao, 2023) Glasses Acoustics Nose No Yes No 11 91.0%
C-Auth (Lim et al., 2023) Glasses Camera Face No Yes Yes 20 96.5%
Zhang et al. (Zhang et al., 2018) Glasses Camera Eyes Yes Yes Yes 30 93.1%
Kawasaki et al. (Kawasaki and Sugiura, 2022) Glasses Optical Blink Yes No No 7 93.6%
EarEcho (Gao et al., 2019) Earables Acoustics Ear Canal No Yes Yes 20 94.5%
ToothSonic (Wang et al., 2022) Earables Acoustics Toothprint Yes No No 25 92.9%-98.9%
WristAcoustic (Huh et al., 2023) Wristband Acoustics Wrist Yes Yes Yes 25 95.0%
2.2. Authentication on Wearable Devices
In recent years, the proliferation of wearable devices, such as smartwatches, smart rings, earphones, and smart glasses, has been notable. These devices, as highlighted by (Ometov et al., 2021), are increasingly integrating into our daily lives, storing sensitive data ranging from SMS messages (Do et al., 2017) to health information (Abdelhamid, 2021). Moreover, their application in critical functions, such as contactless payments (ur Rehman et al., 2022) and authorizing access to laptops (Shahzad and Singh, 2017), underscores the necessity of robust access control systems. Consequently, there has been a significant thrust in research towards developing effective authentication methods for these devices.
2.2.1. Smart Glasses
To enhance the security of smart glasses, a variety of novel authentication methods have been investigated. Traditional PIN-based mechanisms are notably vulnerable to shoulder-surfing attacks (Tari et al., 2006). This issue is exacerbated in smart glasses, where the PIN is entered on a touchpad, making it more visible to bystanders (Islam et al., 2018; Chan et al., 2015). Alternative approaches such as voice-based PIN entry have been explored (Bailey et al., 2014; Yadav et al., 2015; Li et al., 2021). These methods employ a cipher to map the PIN to random digits, obscuring the password from eavesdroppers. However, this technique necessitates mental computations from users, potentially diminishing usability (Bailey et al., 2014). Camera-based authentication methods have also been a focus of research. The Glass OTP (Chan et al., 2015) system utilizes the camera in Google Glasses to scan a QR code on the user’s smartphone for authentication, although this method requires the user to carry a smartphone. C-Auth (Lim et al., 2023), which uses a downward-facing camera in the glasses to capture facial contours for authentication. Zhang et al. (Zhang et al., 2018) developed a continuous authentication system based on eye movement response to visual stimuli, detected by a camera. Another approach involves iris recognition, where internal infrared cameras are used for authentication (Li and Po-Jen, 2017). Behavioral biometrics have also been explored for smart glass authentication. Jagmohan et al. (Jagmohan et al., 2016) demonstrated a gesture-based continuous authentication system. The GlassGuard system (Peng et al., 2017), utilizes behavioral biometrics derived from touch gestures and voice commands for continuous authentication. Kawasaki et al. (Kawasaki and Sugiura, 2022) introduced an authentication method by observing the skin deformation during blinking, employing a photoreflector to measure blinks. Isobe et al. (Isobe and Murao, 2023) introduced an innovative approach using active acoustic sensing. This method involves transmitting acoustic signals through the nose using speakers integrated into the nose pads of the glasses, with the received signals captured by microphones. The acoustic structure of the nose serves as a biometric identifier, though this technique can be affected by the dryness of the nose. Furthermore, this prototype was not tested across different days so the stability of the system is unclear.
2.2.2. Other Wearable Devices
In the domain of smartwatches, researchers have extensively explored biometric-based authentication methods. For instance, Cornelius et al. (Cornelius et al., 2014) investigated the use of on-wrist bioimpedance responses for user authentication. Zhao et al. (Zhao et al., 2020) employed photoplethysmography signals for continuous authentication. Another innovative approach by Watanabe et al. (Watanabe et al., 2021) utilized active acoustic sensing, requiring users to perform four distinct hand poses for authentication. Further, Lee et al. (Lee et al., 2021) proposed a method leveraging vibration responses, measured through accelerometer and gyroscope sensors, to authenticate users. Recent advancements have introduced more sophisticated techniques. WristAcoustic (Huh et al., 2023), constructs a classifier based on three hand poses, utilizing a cue signal emitted from a surface transducer and recorded by a contact microphone. Additionally, WristConduct, (Sehrt et al., 2022), innovatively uses sound waves transmitted through the wrist bones, employing a bone conduction speaker and a laryngophone for authentication purposes.
In addition to on-device authentication methods, researchers have explored approaches that incorporate secondary devices to facilitate user authentication. Nymi band (Nym, [n.d.]) is a wrist-worn wearable device that integrates fingerprint recognition, serving as an authentication mechanism for various devices and applications. Furthermore, ear-based authentication systems such as EarAE (Wu and He, 2023) and EarEcho (Gao et al., 2019) utilize acoustic sensing techniques to capture the unique structure of the ear canal, thereby offering a biometric signature for authentication purposes. ToothSonic (Wang et al., 2022), employs an earable-based system that utilizes toothprint-induced sonic waves for user verification. Similarly, Amesaka et al. (Amesaka et al., 2023) have demonstrated the potential of using sound leakage signals from hearables to capture the acoustic characteristics of the ear canal, auricle, or hand, which can then be employed for authentication. EarDynamic (Wang et al., 2021) leverages acoustic sensing in earables to assess both the static and dynamic motions of the ear canal during speech, providing a novel method for authentication.
2.3. Comparison between SonicID and Other Authentication Methods on Glasses
To better compare our work with prior work of user authentication on glasses, we list all the projects that are most related to SonicID in Tab. 1. As shown in the table, SonicID does not need users’ interaction to perform authentication and achieves comparable balanced accuracy, if not better, to prior work on glasses. This performance is evaluated with a valid number of participants and remains relatively consistent across different remounting sessions and days. In the meantime, SonicID maintains low-power and minimally-obtrusive due to the advantages of acoustic sensors. In summary, SonicID contributes to the field by presenting the first acoustic-based method on smart glasses that scans users’ faces to obtain biometric information for user identification.
3. Facial Biometirc Information Extraction with Active Acoustic Sensing
In this section, we first introduce background on using face-based biometric information for authentication, followed by the core sensing technique and algorithm of SonicID to obtain unique acoustic features of each user. Then we present the deep learning model used in SonicID to authenticate users and the metrics used to evaluate the performance of our system.
3.1. Face-based Biometric Authentication
Biometrics are unqiue physical characteristics that can be used for personal recognition (of Homeland Security, 2024). Face, as one of the most important parts of human body, has been a crucial source from which biometric information is extracted because the geometries of the facial surfaces of different people are different. Extensive research has been carried out to utilize this biometric information from the user’s face for authentication, especially captured by frontal cameras (Beumier and Acheroy, 2000; Bicego et al., 2006). One of the most commonly used authentication methods based on facial biometrics captured by cameras is Apple’s FaceID (Apple, 2024). It delivers impressive performance in authentication and has advantages of being more convenient and secure compared with password-based methods. The core sensing pricinple of SonicID is similar to FaceID. It instruments acoustic sensors on smart glasses to scan the user’s face in 3D with ultrasonic sound waves instead of infrared cameras and uses the obtained biometric information for authentication. The subsequent subsections introduce the principle of operation of SonicID in detail.
3.2. Signal Transmission and Reception
SonicID adopts active acoustic sensing to scan a user’s face as soon as they wear the device. Active acoustic sensing includes an emitter (e.g. speakers) to transmit encoded signals and a receiver (e.g. microphones) to receive the reflected signals. By comparing the received signals and the transmitted signals, one can obtain and analyze the status and change of the objects in the environment that the signals are targeted at.
In SonicID, we choose signals that sweep within a certain frequency range as transmitted signals. Considering SonicID is deployed on smart glasses which is quite close to users when they put it on, we decide to sweep the signals in the ultrasonic frequency bands, i.e. over 18 kHz, to alleviate the impact of the audible sound on users’ daily activities. In order to obtain richer information from users’ face, we put one pair of speaker and microphone on each side of the glasses. The speakers on two sides are designed to transmit encoded signals in different ultrasonic frequency ranges to avoid interference. We pick 18-21 kHz for the speaker on the right side and 21.5-24.5 kHz for the speaker on the left so that they are inaudible to most people and compatible with the audio interfaces on most commodity devices. A sampling rate of 50 kHz is selected to support these two frequency ranges and the signals sweep from the lower frequency boundary (18/21.5 kHz) to the higher frequency boundary (21/24.5 kHz) every 12 ms. Therefore, each sweep contains N = 600 samples ($50kHz\times 12ms$). Fig. 1 (a) demonstrates one of the transmitted signals that sweeps from 18-21 kHz. These signals are commonly used and named as Frequency Modulated Continuous Waves (FMCW) in the field.
In practice, the two speakers emit the encoded signals that sweep in two frequency ranges repeatedly towards the user’s face when the user put on the device. The signals are reflected by the user’s face and captured by the two microphones that are placed next to the speakers. By analyzing the differences between the received signals and transmitted signals, we can obtain unique acoustic features related to this specific user, which will thus be used for user identification. A detailed description of how SonicID generates the acoustic features is presented below.
(a) Transmitted Signal (18-21 kHz)
(b) Received Signal (18-21 kHz)
(c) Cross-correlation Result
(d) Echo Profile
Figure 1. FMCW Signal Transmission, Reception and Processing.
3.3. Echo Profile Calculation
The two microphones capture two channels of audio on two sides of the glasses. An example of the received signal in one channel (18-21 kHz) is shown in Fig. 1 (b). After receiving the signals, we first apply two Butterworth band-pass filters, which are from 18-21 kHz and 21.5-24.5 kHz, on the signals separately to filter out the noises that are outside the frequency range of our interest. By applying two band-pass filters on two channels of signals, we get four channels of filtered signals, two of which represent the signals that travel on the same side of the face while the other two represent the signals that travel across the face from one side to the other side of the face. The acoustic patterns associated to the user’s face that we would like to obtain are called Echo Profiles, which have been demonstrated in several previous papers (Wang et al., 2018; Li et al., 2022; Mahmud et al., 2023; Appear], 2024). According to prior work, we continuously calculate the cross-correlation between the received signals and the transmitted signals using Eq. 1.
(1) $C(n)=Tx(n)*Rx(n)=\sum_{m=0}^{N-1}Tx(m)\cdot Rx(m+n),n\geq 0$
where $Tx(n)$ is the transmitted signal and $Rx(n)$ is the filtered received signal. Since there are four channels of filtered received signals after the band-pass filters, we calculate the cross-correlation between each one of these channels and the corresponding transmitted signal, leading to four channels of acoustic patterns related to different speaker-microphone links. Fig. 1 (c) demonstrates $C(n)$ obtained after the cross-correlation calculation in one channel. The strength of $C(n)$ correlates to the strength of the signals that the system receives at different time. The 0-point is the direct path which means that the signal travels from the speaker directly to the microphone via solid since they are on the same glasses. The positive samples represent the signals that arrive at the system later than the direct path. They usually travel in the air and are reflected by the objects in the surroundings of the system. The negative samples represent those arriving earlier than the direct path and they come from previous frame(s) of the transmitted signals, which are being emitted continuously and repeatedly. Two consecutive samples are 0.02 ms apart ($1\div 50kHz$) and we only show the center 120 samples in Fig. 1 (c).
We then reshape the one-dimensional $C(n)$ by the frame size N = 600 samples to form a two-dimensional array. This array of patterns is defined as Echo Profile, which includes the biometric information of the face of the user who is wearing the device. Fig. 1 (d) shows an example of the echo profile that is properly reshaped from Fig. 1 (c). In echo profiles, the y-axis is the distance, which shows the strength of the signals that is reflected from different distances away from the SonicID system on glasses. Because different users have different facial shape and contours, the echo profile shows the facial scanning result of each user and is distinctive for each user. Considering the average length of people’s faces, we crop 70 samples of echo profiles from -15 samples to 55 samples as the acoustic pattern to authenticate users instead of using the full range of it, which scans the area at a distance of 0 cm to 18.7 cm from the system ($55~{}samples\div 50kHz\times 340m/s\div 2$). We divide the distance by 2 because the signals travel round-way from the system to the face and back to the system. 18.7 cm is the maximum one-way distance from the system that we would like to scan. As stated above, the negative samples are from previous frames and theoretically do not contain much useful information. However, we still include 15 samples from the negative side because our transmitted signals are not infinite in the frequency domain and thus the acoustic patterns diffuse to the negative side to some extent when the cross-correlation is calculated. The cropped echo profiles are fed into a deep learning model to authenticate users, which is introduced in the subsection below.
3.4. Machine Learning Model for Authentication
Though the acoustic data is originally one-dimensional, we reshape the echo profiles to make them two-dimensional images. In order to extract features that are unique to each user from these two-dimensional echo profiles, we decide to adopt a deep learning model based on the ResNet-18 architecture because Convolutional Neural Networks (CNN) are known for being good at distinguishing different classes. Hence, we construct the model using the ResNet-18 architecture as the encoder to extract features from echo profiles and using a fully-connected layer as the decoder. The model serves as a binary classifier which classifies the real user as positive and the attackers as negative. While training the model for a new user, the positive instances come from the data that this user provides when he/she registers on the device and the negative instances come from an existing dataset which was pre-collected from different people.
The input fed into the model is the reshaped and cropped echo profiles. We take 10 600-sample frames as one instance, which corresponds to 0.12 s in time ($10\times 600~{}samples\div 50kHz$). Therefore, the size of each instance as input is $4\times 10\times 70$, where 4 represents the 4 channels of audio data and 70 represents the number of the cropped vertical samples. During testing, the positive instances are still provided by the real user while the negative instances come from the unknown attackers that have not been seen by the model.
In the CNN model, we use an Adam optimizer, a cosine scheduler and set the initial learning rate to be 0.0002. The model is trained to minimize the cross-entropy loss for 50 epochs. Random vertical shift is included and random noise is added as the data augmentation methods to boost the generalizability of the model.
3.5. Evaluation Metrics
To evaluate the performance of our model in SonicID, we adopt three commonly used metrics in prior authentication research, which are True Positive Rate (TPR), False Positive Rate (FPR) and Balanced Accuracy (BAC). The euqations to calculate these three values are as follows.
(2) $True~{}Positive~{}Rate~{}(TPR)=\frac{True~{}Positives}{True~{}Positives+False~{}Negatives}$
(3) $False~{}Positive~{}Rate~{}(FPR)=\frac{False~{}Positives}{False~{}Positives+True~{}Negatives}$
(4) $Balanced~{}Accuracy~{}(BAC)=\frac{TPR+(1-FPR)}{2}$
TPR evaluates the ability of the system to authenticate the real users while FPR evaluates the ability of the system to protect itself from being attacked. BAC gives a balanced evaluation between the two metrics above.
4. Prototype Design and Implementation
This section presents the design and implementation of the hardware prototype. First, we introduce the micro-controller and sensors used in the system. Next, we present how we prototype the system on the glasses form factor.
(a) Hardware Components
(b) Explore Sensor Positions
(c) Prototype
(d) User Wearing Prototype
Figure 2. Hardware Components and Prototype of SonicID.
4.1. MCU and Sensors
Fig. 2 (a) demonstrates the hardware components used in the SonicID system. In order to implement the signal transmission and reception methods introduced in Sec. 3.2, we use the speaker OWR-05049T-38D [^1] https://www.bitfoic.com/detail/owr05049t38d-14578 to emit encoded signals and the microphone ICS-43434222https://invensense.tdk.com/products/ics-43434/ to receive reflected signals. We customize Printed Circuit Boards (PCB) for the speakers and microphones to make them minimally-obtrusive and compatible with different micro-controllers. Teeny 4.1333https://www.pjrc.com/store/teensy41.html is utilized as the micro-controller in the SonicID system to manage the operation of speakers and microphones because of its good performance in processing audio data. Two of the amplifiers MAX98357A [^4] https://www.analog.com/en/products/max98357a.html are used in the system to support as many as 2 speakers and 8 microphones. Specific PCBs are also designed for these amplifiers to make them easily fit the Teensy 4.1 board. Speakers and microphones are connected to Teensy 4.1 via Flexible Printed Circuit (FPC) cables and they communicate with each other with the Inter-IC Sound (I2S) buses. The transmitted signal is pre-programmed into the memory of Teensy 4.1 and the received audio data is stored into the on-board SD card on Teensy 4.1.
4.2. Form Factor
After we select the appropriate components for the SonicID system, the next step is to prototype SonicID on the glasses form factor. In order to validate the feasibility of SonicID, we decide to deploy the system on a pair of commodity glasses. Since we want to scan the user’s face to obtain their biometric information, we would like to point the speakers and microphones downwards, emitting signals directly towards the user’s face. There are only limited space on glasses where we can place the sensors, facing downwards, as shown in Fig. 2 (b): (1) the bridge of the glass frame; (2) the bottom of the glass frame; (3) the hinges of the glasses. If the sensors are placed at position (1) or (2), there is a chance that the sensors may touch the user’s face or block the view of the user because these two positions are right in front of the user’s face. Therefore, we determine to put the sensors at the hinges of the glasses, where there is some space to instrument sensors without letting them touch the user’s face or block the user’s view. Moreover, this sensor position has been leveraged in multiple prior work to place acoustic sensors to track facial expressions (Appear], 2024) and body poses (Mahmud et al., 2023). We can potentially integrate SonicID into their systems to realize activity tracking and user authentication with one set of hardware.
As displayed in Fig. 2 (c), we put one pair of speaker and microphone on each side of the glasses to obtain richer information and the speaker and microphone are close to each other to better capture the reflected signals. The micro-controller Teensy 4.1 is also fixed to the left leg of the glasses with hot glue and tapes to make the entire system wearable. The FPC cables connecting Teensy 4.1 and the speakers and microphones are properly restrained and aligned to the glass frame. The final prototype looks similar to a pair of commercial smart glasses and is comfortable to wear, as shown in Fig. 2 (d).
5. User Study
To evaluate the performance of SonicID, we designed and conducted a user study across several days. The goal of the study is to validate SonicID’s capability of authenticating users and protecting the system from attacks.
5.1. Study Design
We used the prototype described in Sec. 4 to conduct the user study. In addition to the prototype, a laptop (MacBook Pro, 13-inch, Apple M1 chip) was used in the study to show the instruction video to the participants. The study was conducted in an experiment room on campus. When the study started, the participants came in and was introduced the detailed study procedures before signing the consent form. Then they sat in front of the laptop and wore the prototype. Some participants had long hair which sometimes got stuck between the hardware components and the glasses when they wore the device so we provided hair ties for participants to put their hair up if their hair was long.
After the prototype was properly worn, the data collection process started. An instruction video was played on the laptop. The participants were first instructed to stay still for 10 seconds when the system scanned their face continuously and then asked to remount the device within 10 seconds. Remounting meant taking off the prototype and putting it back on. This procedure was designed to simulate real-life wearing experience since users of wearable devices frequently remount the device everyday. This added variance to the collected data. After the prototype was remounted, the participants stayed still for another 10 seconds and repeated the process above. The data collection lasted for 12 minutes for each participant where 36 10-second remounting sessions of data was collected.
The participants were asked to come back on two other days after the first day and collected data two more times with the same procedures because we would like to evaluate the stability of SonicID across different days. After the participants were done with the study on all three days, they filled out a questionnaire to provide their demographic information and feedback on using the prototype and the system.
5.2. Participants
The user study was approved by the Institutional Review Board (IRB) of the home institution of the researchers. We successfully recruited 24 participants (13 females, 9 males and 2 non-binaries) with an average age of 23.8 years old, ranging from 20 years old to 30 years old. For each participant, we collected 18 minutes of data, which were divided in 108 10-second remounting sessions across three days. Hence, 432 minutes of data was collected in total for this user study. Note that the data P11 collected on Day 3 was lost due to the hardware issue so we asked the participant to come back again and redo the study on Day 3.
6. Evaluation
In this section, we present the evaluation results of the user study stated in Sec. 5. We compare the performance of SonicID with different settings. Moreover, we evaluate the usability of the system by analyzing the responses to the questionnaires collected after the study.
6.1. Study Results
6.1.1. Cross-session Performance
In the user study, we collected 36 sessions (6 minutes) of data from each participant on each day. We first evaluated SonicID’s performance across different remounting sessions on the same day. Out of the 36 sessions we collected on Day 1, we considered the first 6 sessions (1 minute) of data as the practice sessions and dumped them while we evaluated the system. Among the remaining 30 sessions, we used the first 24 sessions (4 minutes) of data as positive samples to train the deep learning model described in Sec. 3.4. We would like to scan users’ face while they are staying still. However, they can move and/or blink unconsciously during the study, which might have an impact on the scanning results. Therefore, we divided each 10-second session into multiple instances of 0.12 s, as discussed in Sec. 3.4, and picked top five instances with the smallest movements by comparing the energy change within these instances. Five 0.12-second instances from each session were actually used for training the model. The negative samples used to train the model came from a dataset we collected prior to the user study. To collect this dataset, we ran a pilot study of the same procedures as the formal user study with 17 people, including 7 researchers and 10 people from their networks. Among these people, 7 of them collected data on three days, 2 collected data on two days and the remaining 8 collected data on one day. Each person contributed 5 minutes of data of staying still on each day. All these data in the dataset were used for training the model as negative samples. The same process was conducted to select the most static instances from all the sessions while training.
Figure 3. True Positive Rate for all Participants across Three Days.
Figure 4. False Positive Rate for all Participants across Three Days.
Figure 5. Balanced Accuracy for all Participants across Three Days.
After the model was properly trained for each participant in the user study, the remaining 6 sessions (1 minute) of data from this participant was adopted as positive samples to test the model. The negative samples for testing came from the other 23 participants in the user study. Each of these 23 participants contributed 5 minutes of data to test the model. In this way, these "attackers" during evaluation were not seen by the model in the training process. Static instances were also selected from all sessions for testing. The model made binary classifications to authenticate users and computed the three metrics introduced in Sec. 3.5 to evaluate the results. We trained an individual model for each participant and the average TPR, FPR and BAC across 24 participants is 96.5% (std=9.0%), 4.1% (std=4.6%) and 96.2% (std=5.6%). The detailed metrics for every participant are shown in Fig. 3, Fig 4 and Fig. 5 respectively.
The average FPR of 4.1% indicates that the system rejects false users with a success rate of over 95%. It demonstrates the ability of SonicID to protect the smart glasses from being attacked by bad actors. The average TPR of 96.5% means that the system authenticates the real users successfully with a rate of over 95%. It shows the ability of SonicID to recognize true users of the smart glasses. BAC gives a balanced evaluation of the two metrics above.
6.1.2. Cross-day Performance
The results above validated the performance of SonicID across different remounting sessions on the same day. Then we evaluated the performance across different days, which was meant to validate the stability of SonicID. On Day 2 and Day 3, we also collected 6 minutes of data from each participant using the same procedures as Day 1. To be consistent with Day 1, we also dropped the first 1 minute of data from every participant. Then the static instances picked from the remaining 5 minutes of data were used for testing the cross-day performance of SonicID. We did not retrain the models and directly used the models trained on Day 1. Each participant provided his/her own 5 minutes of data as positive samples to test the model for TPR on Day 2 and Day 3 separately. The negative samples still came from the other 23 participants in the user study to test FPR.
The results showed that the average TPR, FPR and BAC for Day 2 across 24 participants are 91.3% (std=11.9%), 4.5% (std=4.8%) and 93.4% (std=6.2%) while those for Day 3 are 80.7% (std=18.3%), 4.2% (std=5.1%) and 88.3% (std=10.1%). The individual study results for each participant are displayed in Fig. 3, Fig 4 and Fig. 5 respectively. We ran a one-way repeated measures ANOVA test among TPR on three days for all 24 participants and identified a statistically significant difference ($F(2,69)=7.99,p=0.001<0.05$). Specifically, we conducted a repeated measures t-test between TPR on Day 1 and Day 2, and we did not find a quite statistically significant difference ($t(23)=1.75,p=0.09>0.05$). While the same repeated measures t-test was conducted between TPR on Day 1 and Day 3, a statistically significant difference was found ($t(23)=4.00,p=0.0006<0.05$). This indicates that TPR remains consistent on Day 2 but decreases on Day 3, compared to that on Day 1. However, the average TPR is still over 80% even on Day 3. Under the cases when the real users fail to authenticate themselves into the smart glasses, they can adjust the glasses position or remount the device once, and they will likely log into the system successfully. The decrease of TPR can be attributed to the lack of positive training data from the user. One potential solution to improving TPR is to continuously collect the biomentric information of users’ face while they are using the device and update the model accordingly to increase the success rate of authentication. We plan to explore this more in the future.
Similarly, we also ran a one-way repeated measures ANOVA test among FPR on three days for all 24 participants and did not find a statistically significant difference ($F(2,69)=0.39,p=0.68>0.05$). This validates the stability of SonicID to protect the system from attackers across different days. In such authentication systems, it is usually more crucial to keep FPR lower than keeping TPR higher in case of a tradeoff since we generally do not want other people to have easy access to our personal devices.
In Fig. 3 and Fig. 4, we can see that P20 has the lowest TPR, especially on Day 3 while P12 has the highest FPR. By checking the videos of the participants recorded during the study for reference, we figured that P12 and P20 have long hair and bangs which made it harder for them to put the glasses back to the same position every time after remounting. This negatively affected the performance of the system on them. In Sec. 7.4, we discussed several potential methods to improve the system performance and stability in future.
6.2. Impact of Different Amounts of Training Data
(a) True Positive Rate
(b) False Positive Rate
(c) Balanced Accuracy
Figure 6. Average TPR, FPR and BAC across 24 Participants with Different Amounts of Training Data.
In the evaluation results above, we adopted 4 minutes of training data from each participant, which leads to satisfactory performance. However, in practice, the fewer data a new user needs to provide to register on the device, the more user-friendly the system will be. Therefore, we conducted another experiment to explore the impact of different amounts of training data on the system performance. The average TPR, FPR and BAC among 24 participants with different amounts of training data is illustrated in Fig. 6. As shown in the figure, both TPR and FPR decreased when we reduced the amount of data to train the model. From 4 minutes of data to 1 minute of data, TPR drops from 96.5% to 90.3% on Day 1, from 91.3% to 72.3% on Day 2 and from 80.7% to 50.8% on Day 3 respectively. In the meantime, the FPR also decreases from 4.1% to 1.6% on Day 1, from 4.5% to 2.3% on Day 2 and from 4.2% to 1.9% on Day 3. Hence, BAC changes from 96.2% to 94.3% on Day 1, from 93.4% to 85% on Day 2 and from 88.3% to 74.4% on Day 3 correspondingly. These results indicates that the real users have more difficulty authenticating themselves into the device but the system protects the device from unknown users better, with fewer amounts of training data. Since FPR is a more important metric in authentication systems as discussed above, we can just collect 1 minute of training data when a new user registers on the device, leading to a TPR of 90.3% and a FPR of 1.6%. With time passing by, the authentication mechanism might tend to fail more frequently when the registered users want to use the device. However, new training data can be continuously collected every time when the user wears the device and makes some operations. The model will be updated accordingly to make sure that the registered users have no problem getting access to the device.
6.3. Comparison of Two Sides of Sensors
When we designed the prototype for the SonicID system, we instrumented one pair of speaker and microphone on each side of the glasses to collect complete biometric information from users’ face. However, many commercial smart glasses only have the speaker and microphone on one side of the device. As a result, we would like to experiment on how the system performs when only one side of sensors are utilized. As designed in Sec. 3, four speaker-microphone channels are input into the deep learning model when two pairs of speakers and microphones are deployed. If we only want to use on pair of speaker and microphone on either right side or left side of the glasses, we can just input one speaker-microphone channel related to this pair of speaker and microphone into the model. The experiment results are demonstrated in Tab. 2.
Table 2. Average TPR, FPR and BAC across 24 Participants with Different Channels of Signals. Data Format: TPR/FPR/BAC.
Different Channels Day 1 Day 2 Day 3
Both Sides 96.5%/4.1%/96.2% 91.3%/4.5%/93.4% 80.7%/4.2%/88.3%
Right Side 91.8%/3.5%/94.2% 69.7%/4.1%/82.8% 61.1%/3.9%/78.6%
Left Side 81.1%/2.5%/89.3% 38.7%/3.1%/67.8% 21.8%/2.4%/59.7%
As we can see in the table, the system performance gets worse when only one side of sensors are used. We believe this is because deploying the sensors on one side only extract information from the channel of signal that travels on the same side of the face while instrumenting sensors on both sides also obtain information from the channel of signal that travels across users’ face from one side to the other side. This variety of biometric information helps boost the system performance. In addition, one thing we noticed is that right side has much better performance than left side. Given people’s face are mostly symmetric, we speculate that this discrepancy might be caused by the different frequency ranges we use on two sides (right side: 18-21 kHz and left side: 21.5-24.5 kHz). However, we believe that more thorough experiments are needed to verify this assumption. Even though one side of sensors cause the system performance to drop, the performance on Day 1 using right side sensors are still relatively comparable to using both sides sensors. This means that it is possible for SonicID to only include sensors on one side and more training data can be collected automatically when users are wearing the device to improve the system performance on following days.
6.4. Usability
A questionnaire was distributed to participants after the user study to collect their demographic information and feedback on the prototyped wearable device. First, the participants evaluated the overall experience of wearing the device by answering the question "How comfortable is this wearable device to wear around the face? (0 most uncomfortable, 5 most comfortable)". Among 24 participants, the average rating for this question is 3.6 (std=0.7), indicating that the device with the SonicID system is overall comfortable to wear around the face. Next we specifically asked participants to evaluate the weight of the device with the question "How acceptable do you find the weight of our wearable device? (0 most unacceptable, 5 most acceptable)". The average rating is 3.8 (std=0.9), validating that the weight of the device does not cause much burden on users.
In the SonicID system, we utilized two ultrasonic signals in the frequency ranges over 18 kHz. Theoretically, most people are not capable of hearing the sounds emitted from the system. However, due to the limitation of the hardware components, there might be some frequency leakage into the lower frequency range and some users might be able to hear the sound. As a result, we asked the participants to also answer the question "Can you hear the sound emitted from our system?". 18 out of 24 participants answered "No" to this question while the other 6 participants reported "Yes". For those participants who answered "Yes" to this question, a follow-up question "If yes, how comfortable do you feel about the emitted sound? (0 most uncomfortable, 5 most comfortable)" was asked. The average rating for this question among 6 participants is 4.3 (std=0.5). Even though some users might hear the sound from the system, they generally find this sound not bothering them a lot. Since the echo profile we use as an instance to authenticate users is as short as 0.12 s as discussed in Sec. 3.4, we do not expect the sound to trouble the users continuously. Furthermore, it is possible to lower the strength of the transmitted signals to make them less noticeable by users.
Finally, the participants provided some open-ended comments on the wearable device we prototyped. Most participants found this device easy to wear and no discomfort while wearing it. Some participants reported that the glasses was a little imbalanced due to the placement of the micro-controller. Some participants stated that their hair was taken away by the micro-controller when they taken off the device. Other participants believed that the placement of the micro-controller and the wires can be improved to make the prototype more comfortable. These issues can be addressed by balancing the two sides of the glasses and better incorporating the micro-controller into the glasses with some cases or covers. Besides, some participants thought that the glasses was a bit tight for them. This is the problem of the commodity glasses we use instead of the SonicID system itself. However, we can potentially implement several prototypes of the system on glasses of different sizes to tackle this problem.
7. Discussion
7.1. Power Consumption
We wanted to evaluate the power consumption of the SonicID system in operation. For this purpose, we measured the current that flowed through system with a current ranger [^5] https://lowpowerlab.com/guide/currentranger/ when the system was operating. The current was measured at 165.4 mA with a voltage at 3.3 V. This gave us an average power consumption of 545.8 mW. Considering that SonicID only needs an instance of 0.12 s to authenticate users every time, it consumes 19.8 mAs of energy for each authentication trial. If SonicID is deployed on Meta Ray-Ban with a battery capacity of 154 mAh, the authentication system can be activated around 28,000 times in theory, if SonicID is used alone.
7.2. Health Implications
Even though we picked the signals in the ultrasonic frequency ranges to make them inaudible to the users, they may still cause health concerns because the system is close to users’ ears when they wear it. Therefore, we used the NIOSH sound level meter App [^6] https://www.cdc.gov/niosh/topics/noise/app.html to measure the sound level of the signals emitted from the system. We kept the system transmitting signals continuously and placed the microphones of the smartphone with the App running at a distance, where users’ ears would be if they wore the glasses, from the speakers of the SonicID system. The sound level was measured at 65.8 dB. Based on the findings in (Howard et al., 2005), the recommended exposure limit for ultrasonic signals around 20 kHz is 75 dB. This indicates that our SonicID system is safe to wear for the general public given that the authentication process only lasts for a short period of time when users start to operate the device. What is more, the signal strength can be reduced to further restrain its impact on users’ health.
7.3. Applications
SonicID proposes a low-power and minimally-obtrusive solution to user authentication on smart glasses. Meanwhile, the current prototype is relatively cheap (around \$50) with a lot of potentials for further modification and optimization with an even cheaper cost. Therefore, we expect the system to be applied on more wearable devices such as VR headsets, AR glasses, etc. The authentication pipeline can also be utilized in scenarios beyond that when users trying to log into the device. It can be used anytime when you need user identification on wearable devices, including making payments, logging into a social account, changing the settings of the device, etc. Despite the promising application space of SonicID, it is important to keep improving its performance to make sure that the false positive samples are as few as possible.
7.4. Potential Methods to Improve Performance
As evaluated in Sec. 6, although the FPR is consistently below 5% across several days in evaluation, the TPR reduces on Day 2 and Day 3 compared with Day 1 for some participants. We believe this is due to several reasons. Firstly, the positive samples we collected from each participant every day were just from the 5 minutes of data. The small size of the data can lead to fluctuations in the evaluation results in TPR. Secondly, we noticed that TPR is especially low for P20 who has long hair and bangs. The hair can sometimes cover the system and prevent the participant from putting the glasses back to the same position on their head every time they remount the device. In future, we plan to explore several potential methods to improve the performance of the system.
7.4.1. Explore Movement-based Authentication
In Sec. 6, we removed information of blinking and other movements from the data during training and testing to guarantee a valid facial scanning outcome. However, this information can also be taken advantage of to authenticate users. For instance, the blinking information alone can be exploited as a unique biometric pattern because different users blink differently. Furthermore, we can explore other movement-based authentication methods to add another layer of protection. Users of our system can select a specific facial gesture (e.g. smile, tongue out, blink, etc.) as their password to the system. Our system recognizes both their password and their biometric information when performing this password to authenticate users. This kind of two-layer authentication systems make the system more secure and are possible to implement given that there are already some work validating that deploying acoustic sensing on wearables can help track users’ facial movements (Li et al., 2022; Zhang et al., 2023; Xie et al., 2021; Appear], 2024). SonicID can potentially be integrated into these existing systems to realize user authentication without adding extra hardware.
7.4.2. Continuously Collect Training Data
Another potential way to improve the system performance is to continuously collect training data while the user operates the glasses. Currently, SonicID runs a registering process for new users to collect training data. This small amount of training data limits the stability of the system across different days due to the lack of variance, especially for TPR. SonicID can alleviate this problem by continuously collecting the biometric information from the user’s face while they use the smart glasses. The system can scan their face occasionally when they use the device and pick the static instances to train and fine-tune the model. This helps reduce the burden of collecting too much training data before using the authentication system.
7.4.3. Pre-collect More Data in Pilot Study
While we trained the models in the user study, the negative samples came from the 17 people from whom we collected data in the pilot study. If we can recruit more volunteers to collect negative samples in this dataset, the system performance, especially FPR, can be improved since the model will learn the acoustic patterns better from more data. In the meantime, this does not add extra burden of collecting more training data on the new users.
7.5. Limitation and Future Work
SonicID provides a low-power, effortless, and minimally-obtrusive solution to user authentication on smart glasses that works relatively consistently well across different remounting session and different days. However, there are still some limitation of this work.
7.5.1. Impact of Long Hair
The long hair and bangs of users might impact the performance of SonicID and even cover the system. In the study, participants with long hair were asked to put their hair up with a hair tie. Nevertheless, some participants experienced difficulty putting the glasses back to the same wearing position every time and had worse performance in TPR or FPR compared with other participants. In the future, we plan to further upgrade the prototype by placing the MCU and sensors at a better position and covering them with a case so that they are not easily impacted by long hair and bangs.
7.5.2. Amount of Training Data Needed
Sec. 6 evaluates the performance of SonicID with different amounts of training data. With 1 minute of data, SonicID achieves satisfactory performance on Day 1 but the performance, especially TPR decreases on Day 2 and Day 3. One solution to this problem is to collect more training data when the user operates the device continuously in the first couple days of wearing the device as discussed in Sec. 7.4.2. Moreover, when we are able to collect data from more participants in the future, the model can better learn the acoustic patterns and make better authentication.
7.5.3. Adapt to Commodity Smart Glasses
To validate the feasibility of SonicID, we deployed the system on a common pair of glasses and conducted the study. In future, we plan to directly use the built-in speakers and microphones on commercial smart glasses to implement the system so that it is directly applicable and ready to use by customers.
8. Conclusion
In this paper, we present an acoustic-based solution to user identification on smart glasses, called SonicID. It adopts active acoustic sensing to scan the user’s face so that biometric information is obtained from the user to authenticate them. A customized binary classifier based on the ResNet-18 architecture is used to extract unique features related to each user from the acoustic patterns. A user study with 24 participants validated the performance of SonicID across different remounting sessions and days. SonicID authenticates the user within 0.12 second as soon as he/she puts on the device and only consumes 19.8 mAs of energy for each authentication trial. We further explore SonicID’s performance under different settings and discuss the potential applications of SonicID and how it can be deployed on commercial smart glasses in the future.